### v1

# The Prism Hypothesis: Harmonizing Semantic and Pixel Representations via Unified Autoencoding

Abstract
Deep representations across modalities are inherently intertwined. In this paper, we systematically analyze the spectral characteristics of various semantic and pixel encoders. Interestingly, our study uncovers a highly inspiring and rarely explored correspondence between an encoder’s feature spectrum and its functional role: semantic encoders primarily capture low-frequency components that encode abstract meaning, whereas pixel encoders additionally retain high-frequency information that conveys fine-grained detail. This heuristic finding offers a unifying perspective that ties encoder behavior to its underlying spectral structure. We define it as the Prism Hypothesis, where each data modality can be viewed as a projection of the natural world onto a shared feature spectrum, just like the prism. Building on this insight, we propose Unified Autoencoding (UAE), a model that harmonizes semantic structure and pixel details via an innovative frequency-band modulator, enabling their seamless coexistence. Extensive experiments on ImageNet and MSCOCO benchmarks validate that our UAE effectively unifies semantic abstraction and pixel-level fidelity into a single latent space with state-of-the-art performance.
Figure 1: The Prism Hypothesis. Our conceptual “prism” decomposes various natural inputs into spectral components along frequency. Low frequency bands capture global semantics and abstract meaning, while high frequency bands encode local detail and fine visual texture. This motivates our Unified Autoencoding (UAE), which harmonizes semantic and pixel representations within a single latent space.
1 Introduction
Trained on massive corpora, recent foundation models have profoundly reshaped perception and generation systems, generalizing well across diverse downstream tasks [dosovitskiy2021vit, caron2021dino, oquab2023dinov2, radford2021clip]. Yet, early advances in perception and generation evolve along largely separate trajectories. Their objectives are typically distributed across distinct network structures, e.g., employing pretrained semantic encoders [caron2021dino, oquab2023dinov2, radford2021clip], to capture high-level meaning, or pixel encoders [kingma2014vae, razavi2019vqvae2] to compress fine-grained visual detail. While each module excels within its own domain, this fragmentation compels subsequent unification efforts [qu2025tokenflow, huang2025illume+, tang2025unilip] to depend simultaneously on semantic and pixel encoders, forcing networks to reconcile fundamentally heterogeneous representations. The sharp mismatch lowers training efficiency and induces representational conflicts, with these incompatible features often interfering rather than complementing one another.
This fragmentation uncovers a deep-seated tension between abstraction and fidelity, which is a driving force in shaping subsequent foundation models. To alleviate it, recent studies [zheng2025rae, yue2025uniflow, shi2025svg] attempt to transfer semantic encoders into visual generation domains alongside strong pixel decoders. This strategy substantially accelerates convergence and improves semantic correspondence, yet remains limited in recovering fine-grained visual details. In parallel, another research direction seeks to endow pixel encoders with semantic awareness through text supervision [xu2022groupvit, yi2023simseg, wu2024imagetext], semantic encoder distillation [chen2023zeroseg, huang2023g2sd], and hierarchical feature integration [wu2024imagetext]. While these efforts move toward unifying understanding and generation representations in a single module, they often achieve coexistence through trade-offs rather than genuine integration.
At the core of this development lies a fundamental question: How is information about the world represented such that multimodal inputs share a common semantic meaning while preserving their native granularity of detail?
Impressively, we empirically observe that pre-trained semantic features, whether derived from text or vision, tend to reside at the coarse end of the decoupled feature spectrum, primarily capturing low-frequency structures such as categories, attributes, and relations. In contrast, pre-trained pixel-level features extend toward the finer end of the spectrum, representing higher-frequency components that convey intricate appearance and geometric detail. Strikingly, these complementary representations can be harmoniously integrated within a unified encoder, extremely aligning well with the spectral arrangement and fostering a progressive synergy from semantic perception to detailed reconstruction. We posit this as the Prism Hypothesis in Figure 1: the real-world inputs project onto a continuous shared feature spectrum, and what we perceive as different modalities are distinct slices of this underlying continuum.
From this perspective, we introduce Unified Autoencoding (UAE), a tokenizer that learns a shared latent space and harmonizes semantic structure and pixel-level detail. Specifically, UAE features an innovative frequency-band modulator that factorizes real-world content into a fundamental semantic band and residual pixel bands with controllable fine granularity. This design is not only grounded in comprehensive empirical evidence but also validated across diverse reconstruction and perception tasks. With its compact yet semantically expressive representations, UAE outperforms concurrent RAE [zheng2025rae], SVG [shi2025svg], and UniFlow [yue2025uniflow] across rFID, PSNR, gFID and Accuracy metrics on ImageNet, demonstrating that its learned latent space is both semantically representative and pixel-faithful.
Our work makes the following contributions.
  1. 1.
    We introduce the Prism Hypothesis, a unified spectral perspective connecting natural inputs via a shared fundamental band and distinct modality-specific bands, supported by extensive empirical observations.
  2. 2.
    We present Unified Autoencoding, a simple yet effective tokenizer that combines a frequency band modulator and seamlessly aligns with existing diffusion transformers.
  3. 3.
    Our UAE demonstrates optimal reconstruction quality on ImageNet and MS-COCO together with diffusion-friendly latents, and extensive ablations prove the effectiveness of the proposed frequency-band factorization.
2 Related Work
Unified Tokenizers and Unified Representations.
Unifying the representations between pixel and semantic embeddings has become a central objective for existing foundation models. Joint embedding approaches align images and text in a shared representation and enable strong zero-shot transfer [radford2021clip, jia2021align], and have been extended to many modalities, e.g., audio, depth, thermal, and inertial signals [girdhar2023imagebind]. In parallel, modality-agnostic backbones aim to build a unified architecture that can process diverse input modalities and generate task-specific outputs through learned queries or a shared token representation [jaegle2021perceiverio, lu2022unifiedio, lu2024unifiedio2, wang2023beit3, baevski2022data2vec].
On the tokenizer side, discrete codebook methods have demonstrated that the design and granularity of visual tokens play a crucial role in determining how effectively a single sequence model can adapt to vision tasks [esser2021taming, yu2024magvitv2]. Recent work goes further and seeks tokenizers that support both understanding and generation at the same time. OmniTokenizer [wang2024omnitokenizer] learns a joint image video tokenizer with a spatial-temporal decoupled transformer and reports strong reconstruction and synthesis across both domains. Very recent studies deepen this trend. UniFlow [yue2025uniflow] proposes a unified pixel flow tokenizer that adapts a pretrained encoder through layer-wise self-distillation and employs a lightweight patch-wise flow decoder, explicitly targeting the long-standing tension between semantic abstraction and pixel-faithful reconstruction. Two concurrent works remove the traditional variational bottleneck entirely and build unified representation latents for diffusion transformers. Diffusion Transformers with Representation Autoencoders (RAE) [zheng2025rae] replace the usual reconstruction-only encoder with a pretrained representation encoder, e.g., DINO or SigLIP, and a trained decoder, arguing that semantically rich latents accelerate convergence and improve generative fidelity. SVG [shi2025svg] trains a diffusion model on DINO features with a small residual branch for details, reporting faster training, few-step sampling, and improved quality.
Our UAE aligns with this shift. It serves as a unified tokenizer that decouples continuous latent features explicitly corresponding to the underlying spectral structure by factorizing real-world contents into a low-frequency base and residual high-frequency bands. It anchors semantics in the core representation while relegating fine-grained details to residuals for progressive reconstruction.
Frequency and Multi-Resolution Modeling.
Classical image synthesis adopted pyramids and wavelets to separate structure by scale, enabling coarse-to-fine generation and targeted refinement of detail. A typical example is the Laplacian pyramid of adversarial networks [denton2015lapgan], which trains a generator per level and synthesizes images by successively adding higher-frequency residuals. Subsequent analyses of neural networks from a spectral perspective showed that standard architectures prioritize low frequencies and learn higher frequencies later, a phenomenon known as spectral bias [rahaman2019spectral]. Two lines of work respond to this bias. The first uses input or architecture design to improve access to high frequencies, e.g., Fourier feature mappings and periodic activations that help multi-layer perceptrons represent fine detail [tancik2020fourier, sitzmann2020siren]. The second introduces frequency-aware objectives and signal processing choices, such as focal frequency loss to emphasize hard frequencies and alias-free synthesis to avoid spurious high-frequency artifacts [jiang2021ffl, karras2021aliasfree].
Modern generative models retain this multi-resolution view. Cascaded diffusion trains models at increasing resolutions, allowing each stage to learn the appropriate frequency band and its own error distribution [ho2022cascaded]. Variants construct explicit feature pyramids or hierarchical patch schedules, enabling efficient diffusion on large images and video while preserving high-frequency detail [fan2023frido, skorokhodov2024hpdm]. Recent latent diffusion designs introduce cross-magnification spaces or zoomable pyramids that share information across scales and enable large-image reconstruction without retraining [yellapragada2025zoomldm]. In autoregressive models, VAR [tian2024var] casts generation as predicting the next scale or resolution, showing strong ImageNet results and clean scaling trends. This explicit progression from global layout to fine detail emerges as a viable alternative to diffusion. Building on this idea, Next Visual Granularity (NVG) generation [wang2025nvg] produces sequences at a fixed resolution but with progressively finer token granularity, surpassing prior VAR baselines while maintaining structured control over detail. Similarly, NFIG [huang2025nfig] performs discrete next-frequency prediction, demonstrating strong generative performance on the ImageNet benchmark. Our approach aligns with these trends but focuses on unified representation rather than the generator’s training schedule.
3 Methodology
Figure 2: Frequency energy distribution. Normalized energy $e(k)$ across frequency bands for diverse tokenizers. DINOv2 and CLIP focus on low-frequency (semantic) content, while SD-VAE retains more high-frequency energy, capturing finer details.
Figure 3: Overall architecture of our proposed Unified Autoencoding (UAE). The input image is separately encoded by both a pretrained Semantic Encoder (e.g., DINOv2) and the trainable Unified Encoder. The unified encoder is initialized from the semantic encoder and optimized under two complementary objectives: a semantic-wise loss that aligns low-frequency components decomposed from the semantic encoder’s representations, and a pixel-wise reconstruction loss that enforces visual fidelity via the Pixel Decoder by adaptively dilating the high-frequency components. The decoder employs spectral transform blocks to refine residual-frequency content and produce the reconstructed image. This joint optimization harmonizes semantic structure and pixel detail within a single latent space.
3.1 Preliminary Findings
The Prism Hypothesis. Natural inputs are regarded as projections of a common real-world signal onto a shared frequency spectrum. Semantic encoders emphasize a compact low-band that carries categories, attributes, and relations, while pixel encoders observe the same fundamental base together with higher bands that encode edges, textures, and fine appearance. The hypothesis indicates that cross-modal alignment depends primarily on the shared low band.
Formalization. Here, $\mathcal{F}$ and $\mathcal{F}^{\dagger}$ denote two-dimensional discrete Fourier transform and its inverse. For an image $I\in[0,1]^{3\times H\times W}$ and a smooth radial mask $\mathbf{M}^{\mathrm{LP}}_{\rho}$ that passes frequencies within normalized radius $\rho\in(0,1]$,
$I^{\mathrm{LP}}_{\rho}\;=\;\mathcal{F}^{\dagger}\!\bigl(\mathbf{M}^{\mathrm{LP}}_{\rho}\odot\mathcal{F}(I)\bigr),$ (1)
$I^{\mathrm{HP}}_{\rho}\;=\;\mathcal{F}^{\dagger}\!\bigl(\mathbf{M}^{\mathrm{HP}}_{\rho}\odot\mathcal{F}(I)\bigr),$ (2)
where $\mathbf{M}^{\mathrm{HP}}_{\rho}$ is complementary to $\mathbf{M}^{\mathrm{LP}}_{\rho}$, and both masks use cosine transitions to limit ringing artifacts. All filtering is performed in linear space prior to any model-specific normalization. With a frozen vision-language encoder $E$, we compute cosine similarities between text embeddings and image embeddings from $I$, $I^{\mathrm{LP}}_{\rho}$, and $I^{\mathrm{HP}}_{\rho}$. If semantic alignment is carried by the shared base, then the retrieval score $\mathcal{R}$ satisfies $\mathcal{R}_{\mathrm{LP}}(\rho)\ \text{is nondecreasing in}\ \rho,$ $\mathcal{R}_{\mathrm{HP}}(\rho)\ \text{is nonincreasing in}\ \rho$, and $\mathcal{R}_{\mathrm{HP}}(\rho)$ approaches chance once the shared base is removed.
Empirical Verification. Here, we validate the Prism Hypothesis via two complementary analyses: frequency energy decomposition and retrieval robustness under spectral filtering. Exp1: In Figure 2, we first measure the normalized frequency energy distribution $e(k)$ across encoders by averaging the magnitude spectrum of each latent feature channel. This metric quantifies how much representational energy is allocated to each frequency band, revealing that semantic encoders such as DINOv2 and CLIP concentrate most of their energy in the low-frequency region (k=0k{=}0), whereas pixel-oriented models like SD-VAE retain stronger mid- and high-frequency components that capture fine details. Exp2: To assess how these spectral tendencies affect semantic alignment, we conduct a controlled retrieval experiment in Figure 4, using text–image retrieval recall (R@5) as the evaluation metric. Images are progressively filtered with radial low-pass or high-pass masks of increasing cutoff (fraction of Nyquist), while text features remain unchanged. R@5 remains stable under low-pass filtering until the lowest bands are removed, but drops sharply under high-pass filtering, approaching random chance. Together, these results confirm that cross-modal alignment and semantic consistency primarily reside in the shared low-frequency base of the latent spectrum, while higher bands encode modality-specific, fine-grained visual detail.
3.2 Unified AutoEncoder
Using DINOv2 as an example, we initialize our UAE from the pretrained model and remove the register tokens, leaving $C$ channel patch tokens. Input images are center-cropped and resized to the encoder input size, i.e., $224$. The resulting tokens are then reshaped into a latent grid $\mathbf{z}\in\mathbb{R}^{B\times C\times H\times W}$ for subsequent frequency-domain processing.
Figure 4: Retrieval results via frequency filtering. Text–Image retrieval (R@5) remains stable under low-pass filtering but degrades sharply under high-pass filtering, confirming that semantic alignment primarily resides in low-frequency components.
3.3 Residual Split Flow
Given a latent grid $\mathbf{z}\in\mathbb{R}^{B\times C\times H\times W}$, we decompose $\mathbf{z}$ into $K$ frequency bands $\mathbf{z_{f}}\in\mathbb{R}^{B\times K\times C\times H\times W}$ using an FFT Band Projector followed by an Iterative Split procedure. Inspired by the coupling mechanism in flow-based models such as Glow [kingma2018glow], the decomposition is performed in the frequency domain via FFT-based projection, allowing each band to capture a distinct spectral component while maintaining invertibility and spatial consistency.
FFT Band Projector. To isolate frequency components, we first apply a two-dimensional discrete Fourier transform to the latent grid, obtaining complex-valued spectral coefficients. Each frequency band is defined by a smooth radial mask $\mathbf{M}_{k}$, which partitions the Fourier magnitude into $K$ concentric rings. These masks overlap slightly to ensure continuity across spectral boundaries and to prevent ringing artifacts when reconstructed. For each band $k$, they are transformed back to the spatial domain via an inverse FFT, yielding a spatially coherent band-limited feature.
Iterative Split. With the latent representation $\mathbf{z_{f}}^{(0)}=\mathbf{z}$, and residual of frequency $\mathbf{r}^{(0)}=\mathbf{z}$, we iteratively decompose the original feature into $K$ frequency bands as:
$\mathbf{z_{f}}^{(k)}=\mathcal{P}_{k}\!\left(\mathbf{r}^{(k)}\right),$ (3)
where the residual term $\mathbf{r}$ is defined as:
$\mathbf{r}^{(k+1)}=\mathbf{r}^{(k)}-\mathcal{P}_{k}\!\left(\mathbf{r}^{(k)}\right).$ (4)
Here, $\mathcal{P}_{k}$ denotes the FFTBandProejector corresponding to the $k$-th radial mask $M_{k}$. At the first step (k=0k{=}0), the low-frequency base $\mathbf{b}^{(0)}$ is directly extracted without residual subtraction. For subsequent steps (k>0k{>}0), the per-band latent $\mathbf{z_{f}}^{(k)}$ is progressively calculated by removing each extracted band. After $K$ iterations, the decomposition yields the full set $\{\mathbf{z_{f}}^{(k)}\}_{k=0}^{K-1}$ as multi-band latent representations.
This residual decomposition encourages spectral disentanglement in one latent space. Low-frequency bands encode global semantics and smooth structures, while higher-frequency bands capture localized edges and fine details.
3.4 Frequency Band Modulator
After the latent representation $\mathbf{z}$ is decomposed into multiple frequency bands, the bands are processed by the frequency modulation module, mapped back to the spatial domain, and then passed through the ViT decoder to reconstruct the RGB image. During training, the multi-band latent is first processed by a noise injection module that perturbs only the high-frequency bands to improve robustness. The resulting latent is then passed through a spectral transform module, which maps it back to the spatial domain before being decoded by the ViT-based pixel decoder.
Noise Injection.
To further boost the capacity of the decoder, we perform noise corruption during the training. Specifically, we build a binary condition $\mathbf{m}\in\{0,1\}^{B\times K}$ indicating which bands are kept from corruption. For each band tensor $\mathbf{b}^{(k)}$ we form noise injection through:
$\widetilde{\mathbf{b}}^{(k)}=\mathbf{m}_{:,k}\odot\mathbf{b}^{(k)}+\overline{\mathbf{m}}_{:,k}\odot\mathcal{N}\!\left(\mathbf{0},\sigma^{2}\mathbf{I}\right),$ (5)
where $\sigma$ is drawn from $[0,1]$.
Spectral Transform.
As shown in Figure 3, we concatenate processed bands along channels, obtain $\mathbf{s}=\operatorname{Concat}_{\text{ch}}\left(\{\widehat{\mathbf{b}}^{(k)}\}_{k=0}^{K-1}\right)\in\mathbb{R}^{B\times KC\times H\times W}$, and pass $\mathbf{s}$ through a two layer convolutional block with SiLU to predict a residual $\Delta\in\mathbb{R}^{B\times C\times H\times W}$. Note that the final fused latent can be formulated as follows:
$\mathbf{q}=\Delta+\sum_{k=0}^{K-1}\widehat{\mathbf{b}}^{(k)}\in\mathbb{R}^{B\times C\times H\times W}.$ (6)
Notably, this tensor $\mathbf{q}$ maintains the same shape as the encoder output, regardless of how many bands are retained, and is the only tensor consumed by the decoder.
3.5 Semantic-wise Loss
To retain the semantic priors of the pretrained teacher while expanding frequency coverage, we apply a semantic-wise loss between the unified encoder and the frozen semantic encoder on the lowest $K_{\text{base}}$ frequency bands only. Let $\mathbf{f}_{s}$ and $\mathbf{f}_{u}$ denote the features from the pretrained-semantic encoders and trainable-unified encoders, respectively. After frequency decomposition, we obtain band-wise representations $\{\mathbf{f}^{k}_{s}\}_{k=0}^{K-1}$ and $\{\mathbf{f}^{k}_{u}\}_{k=0}^{K-1}$. The semantic alignment is enforced only for the first $K_{\text{base}}$ (low-frequency) bands that encode global structure and object semantics:
$\mathcal{L}_{\text{sem}}=\frac{\sum_{k=0}^{K_{\text{base}}-1}\|\mathbf{f}^{k}_{u}-\mathbf{f}^{k}_{s}\|_{2}^{2}}{K_{\text{base}}}.$ (7)
This restricted alignment ensures that UAE inherits the semantic layout and category-level organization from the teacher in the low-frequency domain, while leaving the remaining high-frequency bands unconstrained for learning modality-specific, pixel-level detail. In practice, this selective supervision stabilizes training and encourages the unified encoder to harmonize semantic and pixel representations without collapsing into purely pixel features. Noted that in our experiments, we set $K_{\text{base}}=1$ by default.
4 Experiments
4.1 Implementation Details
Dataset. Following the standard settings, we train our UAE on ImageNet-1K train set and evaluate it on ImageNet-1K val set and MS-COCO 2017 dataset for fair comparisons.
Training Settings. We train UAE with DINOv2-B and DINOv2-L respectively in 256x256 resolution. The training processes are decomposed into three stages. In stage 1, the semantic encoder is frozen, where the decoder is trained by reconstruction loss. In stage 2, we unfreeze the encoder and finetune the full model with semantic-wise loss and reconstruction loss. In stage 3, we finetune the full model end-to-end with noise injection and GAN loss, where the GAN loss setting follows the representative RAE [zheng2025rae].
Table 1: Reconstruction quality on ImageNet-1K and MS-COCO 2017 (256x256). Note that our proposed UAE achieves state-of-the-art reconstruction among unified tokenizers and remains competitive with strong autoencoders such as Flux-VAE and SD3-VAE. Under identical DINOv2 encoders, UAE significantly outperforms the RAE baseline in both PSNR and SSIM while reducing rFID by over 90%. When scaled to DINOv2-L, UAE attains the best overall perceptual quality (rFID=0.16) and fidelity (PSNR=33.08, SSIM=0.94), demonstrating the effectiveness of frequency-aware factorization in preserving both semantic and fine-grained visual detail.
Method Type Ratio ImageNet-1K MS-COCO 2017
[…]
SD-VAE [rombach2021highresolution] Continuous 8 25.68 0.72 0.75 25.43 0.73 0.76
SD-VAE-EMA [rombach2021highresolution] Continuous 8 24.99 0.71 0.63 24.76 0.72 0.57
SD3-VAE [lopez2025sd3] Continuous 8 29.58 0.86 0.21 29.50 0.87 0.19
Flux-VAE [labs2025flux1kontextflowmatching] Continuous 8 31.07 0.89 0.17 30.99 0.90 0.19
TokenFlow [qu2025tokenflow] Discrete 16 21.41 0.69 1.37 - - -
DualViTok [huang2025illume+] Discrete 16 22.53 0.74 1.37 - - -
EMU2 [Emu2] Continuous 14 13.49 0.42 3.27 - - -
UniLIP [tang2025unilip] Continuous 32 22.99 0.75 0.79 - - -
RAE (DINOv2-B) Continuous 14 18.05 0.5 2.04 18.37 0.49 2.56
[…]
UniFlow (DINOv2-L) Continuous 14 32.32 0.91 0.17 32.29 0.90 0.18
[…]
4.2 Visual Reconstruction
Quantitative Evaluation. In Table 1, We quantify reconstruction quality at $256\times 256$ on ImageNet-1K and MS-COCO 2017. The compared baselines include widely-used generative tokenizers and variational decoders. For a fair comparison, we report the evaluation results for our UAE and the corresponding baseline that follows the same configurations of DINOv2-base. Here, we report PSNR and SSIM for fidelity, and rFID for perceptual quality.
Figure 5: Qualitative comparison of reconstruction fidelity across autoencoding paradigms. We visualize reconstructed samples from representative methods, including SD-VAE [rombach2022ldm], RAE [zheng2025rae], and our proposed UAE. Each row corresponds to reconstructions from a fixed source set spanning text, human, object, and artistic domains. UAE produces the most consistent and semantically faithful reconstructions, preserving both high-frequency details (e.g., texture and edge sharpness) and global structure (e.g., layout and color harmony), while reducing the blurring and semantic drift observed in SD-VAE and RAE. (The detail comparisons are denoted in the yellow boxes.)
Notably, our UAE delivers state-of-the-art reconstruction quality among unified tokenizers. On ImageNet-1K, UAE improves over the RAE baseline from $18.05$ to $29.65$ in PSNR and from $0.50$ to $0.88$ in SSIM, while reducing FID from $2.04$ to $0.19$. On MS-COCO, UAE raises PSNR from $18.37$ to $29.23$ and SSIM from $0.49$ to $0.89$, with FID decreasing from $2.56$ to $0.18$. These notable gains validate that moving masking out of the decoder and factorizing latents into a low-frequency base and residual bands preserve fine detail while maintaining semantic structure.
Beyond unified tokenizers, UAE is competitive with the best generative-only tokenizers. On ImageNet-1K, UAE attains an FID of $0.19$, matching SD three VAE and approaching Flux VAE, while maintaining high PSNR and SSIM. A similar pattern holds on MS-COCO, where UAE equals or surpasses the strongest baselines in perceptual quality.
Qualitative Evaluation. Figure 5 compares source images with the resulting reconstructions from SD-VAE, RAE, and UAE. Note that our UAE can well preserve straight edges, fine textures, and small text, e.g., street signs and printed documents. The improvement is consistent across natural photos and illustrations, as shown by the montage on page one of the uploaded figure. The detailed comparisons are shown in the yellow boxes in the images.
4.3 Generative Modeling
Table 2: Class-conditional generation performance on ImageNet (256x256). We compare our proposed UAE with recent diffusion and autoregressive models using standard metrics. Note that the UAE performs generation in a causal manner, progressing from low- to high-frequency bands in the latent space.
[…]
DiT-XL/2 [peebles2023dit] 2.27 271.3 0.83 0.57
VAR [tian2024var] 1.73 350.2 0.82 0.6
UniFlow [yue2025uniflow] 2.45 228.0 - -
RAE [zheng2025rae] 1.51 242.9 0.79 0.63
[…]
To further assess the effectiveness of the proposed UAE, we conduct class-conditional image generation experiments on ImageNet-1K at a resolution of $256\times 256$.
We conduct all generative experiments in the multi-band latent space. All experimental settings follow those used in RAE [zheng2025rae] to ensure a fair comparison. As shown in Table 2, our UAE attains a gFID of 1.68 and an IS of 301.6, achieving performance on par with existing state-of-the-art generative models. This suggests that the unified frequency-based representation enables the generative model to progressively capture both global structure and fine-grained details in a coherent manner, keeping highly generative while preserving strong semantic quality. Overall, the UAE latent space, constructed through explicit low- and high-frequency decomposition, provides an effective and diffusion-friendly foundation for large-scale visual generation.
4.4 Semantic Understanding
To evaluate the learned semantic discriminability, we perform linear probing on ImageNet-1K following the standard protocol in [yue2025uniflow]. All encoder weights are frozen, and a single linear classifier is trained on top of the global embeddings for 30 epochs using SGD with a cosine learning rate schedule and a base rate of 0.1. Consistent augmentations are applied across methods for a fair comparison.
Table 3: Linear probing on ImageNet-1K. UAE achieves $83.0\%$ top-1 accuracy with a ViT-B backbone, matching RAE and surpassing larger ViT-L models such as MAE, MAGE, and UniFlow. This shows that UAE’s unified latent space preserves strong semantic discriminability with compact model size.
Methods Size $\text{ACC}\!\uparrow$
VFMTok [zheng2025vision] ViT-L 69.4
BEiT [beit] ViT-L 73.5
MAE [he2022mae] ViT-L 75.8
MAGE [li2023mage] ViT-L 78.9
UniFlow [yue2025uniflow] ViT-L 82.6
RAE [zheng2025rae] ViT-B 83.0
[…]
Figure 6: t-SNE visualization of semantic embeddings. We compare the feature distributions from the DINOv2 encoder (left) and the band-0 (low-frequency) component of UAE (right). The two plots exhibit similar global structures and class separability, indicating that UAE effectively preserves the semantic organization of the original encoder while introducing a unified latent space that remains compatible with frequency-based factorization.
As shown in Table 3, our UAE achieves a top-1 accuracy of 83.0% based on a ViT-B backbone, outperforming previous self-supervised and tokenizer-based approaches. Compared to VFMTok (69.4%) and BEiT (73.5%), our UAE exhibits a substantial improvement, demonstrating that the unified latent space preserves stronger semantic alignment. Even against large-scale models such as MAGE (78.9%) and UniFlow (82.6%), UAE maintains competitive classification accuracy despite its smaller backbone.
To further visualize this property, Figure 6 shows t-SNE comparisons between the DINOv2 encoder and the band-0 (low-frequency) component of UAE. The two plots share highly similar global structures and class separability, confirming that UAE’s low-frequency representations retain the original semantic organization of DINOv2, while enabling a unified latent space compatible with frequency-based factorization. This qualitative visualization supports our design goal of preserving global semantics while enriching fine-grained visual detail in higher-frequency bands.
5 Ablation Study
5.1 The Effectiveness of the UAE Modules
Table 4: Configurations of UAE modules. Progressively integrating various strategies improves both fidelity and perceptual realism. The BandProjector enhances structure recovery, tuning pre-trained encoders largely improves pixel-level fidelity, and noise injection stabilizes training for the best reconstruction quality.
[…]
Baseline 15.27 0.45 22.98
+ BandProjector 22.13 0.71 15.22
+ Tuning Encoder 29.02 0.88 0.21
+ Noise Injection 29.65 0.88 0.19
In Table 4, we evaluate the influence of each core configuration in our proposed UAE on ImageNet-1K. Starting from a baseline that directly reconstructs images from the frozen DINOv2-B encoder, we progressively add the BandProjector, Encoder Tuning, and Noise Injection strategies.
Note that these components introduce solid and progressive improvements for visual fidelity. The BandProjector enhances structural recovery by decomposing latent signals across frequency bands, increasing PSNR from 15.27 to 22.13 and SSIM from 0.45 to 0.71. Tuning the encoder further aligns the latent representation with the reconstruction objective, raising PSNR to 27.02 and reducing rFID from 15.22 to 7.81. Finally, the noise injection strategy yields a large gain, achieving PSNR 29.65, SSIM 0.88, and rFID 0.19, indicating impressive perceptual fidelity.
Together, these results confirm that frequency decomposition, semantic alignment, and controlled noise regularization jointly enable UAE to harmonize semantic structure and fine visual detail in one unified representation space.
5.2 The Configurations of Band Splits
Table 5: Configurations different band splits. UAE maintains consistent and stable reconstruction and semantic performance across different numbers of frequency bands, demonstrating the robustness of the proposed frequency factorization design.
[…]
10 29.65 0.88 0.19 83.0
8 29.37 0.88 0.19 83.0
6 29.41 0.88 0.19 83.0
4 29.52 0.88 0.19 83.0
2 29.55 0.88 0.19 83.0
We further examine how the number of frequency bands affects reconstruction fidelity and semantic consistency. In Table 5, varying the number of bands from 2 to 10 produces nearly identical performance across all metrics: PSNR stays around $29$, SSIM remains at $0.88$, rFID is consistently $0.19$, and linear probing accuracy (ACC) stays fixed at $83.0\%$.
These results demonstrate that UAE’s frequency factorization is highly robust to band granularity. Its decomposition reliably preserves both low-frequency and high-frequency information even when the spectrum is split coarsely. The minimal variation further indicates that most of the frequency energy is captured by the base and the first few residual bands, confirming that the unified representation does not depend on fine-grained spectral partitioning.
In practice, using fewer bands slightly reduces computation without harming reconstruction quality or downstream performance. With more multi-frequency latents, our UAE also naturally supports iterative image generation.
5.3 The Semantic Comparisons in DINOv2
Table 6: Semantic comparisons in DINOv2. Comparison of classification accuracy when probing different representations derived from DINOv2. Band0 denotes using only the lowest-frequency band for linear probing, whereas Concat refers to probing with the concatenation of all frequency-band representations.
Methods DINOv2 $\textbf{Band}_{\textbf{0}}$ Concat
Accuracy(%) 83.0 83.3 83.0
As shown in Table 6, using only the lowest-frequency component (Band0) achieves slightly higher accuracy (83.3%) than both the original DINOv2 features and the concatenated multi-band representation (83.0%). This indicates that the low-frequency band effectively retains the global semantic structure of DINOv2, which dominates discriminative performance in classification.
6 Conclusion
We propose the Prism Hypothesis, which views diverse natural inputs as projections of a shared spectrum composed of a compact low-frequency semantic component and residual higher-frequency detail. Preliminary experiments about frequency-band distributions of various visual and textual encoders strongly enlighten this connection between feature spectra and representational function. Hence, we launch Unified AutoEncoding (UAE), which harmonizes semantic and pixel information within a single latent space via a hierarchical frequency-band modulator. Extensive experiments confirm that it delivers greater generative capability over existing unified tokenizers, e.g., RAE, SVG, and UniFlow, meanwhile preserving reconstruction quality on par with top-tier models like Flux-VAE. We position the UAE as a promising and practical route towards unified tokenizers for understanding and generation. .